OpenStudy (anonymous):

What is one of the solutions to the following system of equations? y2 + x2 = 65 y + x = 7 (2 points) (-8, 1) (1, 6) (6, 1) (9, -2)

OpenStudy (usukidoll):

plug them in and see if they equal to 65 and 7

OpenStudy (usukidoll):

I see -8,1 satisfying the first and 9,2 satisfying the second...

OpenStudy (anonymous):

didnt even think of it that way thank you

OpenStudy (raden):

y + x = 7 1 + (-8) = 7 -7 = 7 nah, this is not working :)

OpenStudy (usukidoll):

I said the first ^^

OpenStudy (usukidoll):

1^2 + (-8)^2 = 65 1+64 = 65 65 =65

OpenStudy (raden):

maybe there is a typo, i think. it is (8,-1)

OpenStudy (usukidoll):

hmmmmm that's odd I can see the second, third, and fourth satisfying the second one

OpenStudy (usukidoll):

could be a typo.. 8,-1 does work on both equations